OpenStudy (juan337492):

One number is 7 more than another. The difference between their squares is 231. What are the numbers?

OpenStudy (jiteshmeghwal9):

Let the smaller number be x & larger number be y then y= x+7 .......(1) \[y^2-x^2=231......(2)\] solve this two equation for x & y

OpenStudy (juan337492):

so x(x+7)=231

OpenStudy (otherworldly):

when it says the difference between their squares is 231 it means that when u square the numbers after u find them the difference between them is 231 it never said what the sum is for their squares

OpenStudy (jiteshmeghwal9):

The difference between the squares of number means y^2-x^2

OpenStudy (otherworldly):

=231.....